use the substitution method to solve the linear system. x-y= -5 and x+4=16
Answer by stanbon(75887)

use the substitution method to solve the linear system.
x-y = -5
x+4 = 16
---------------
x = y-5
Substitute for "x" and solve for "y":
y-5+4 = 16
y - 1 = 16
y = 17
---------
x = y-5
x = 17-5 = 12
====================
Cheers,
Stan H.
====================